OpenStudy (salwan):

find the slope x intercept y intercept x=3y+2

OpenStudy (stefrheart):

make into slope intercept form

OpenStudy (anonymous):

To find your x, and y-intercepts, you need to isolate each variable by itself (x is already like that).\[x = 3y + 2\]\[3y=-x+2\]To get the y by itself you subtract 3y to get -3y on the left side and then you subtract the x to make it go on the right side as -x. Once you have that already, to find the x-intercept, what must the y value be? The line crossing the x-intercept must have a value of 0 for y, and vice-versa with the y-intercept. So find the x intercept by subbing in 0 for the variable y, and the same for the other.\[x = 3(0) + 2\]\[-3y = -(0) +2\]When you have done that you will find out that the numbers would equal 0 since anything multiplied by zero is zero. Divide by -3 to get the POSITIVE value for y since a negative variable is not wanted.\[x = 2\]\[y = -\frac{ 2 }{ 3 }\]When you have your x and y-intercepts, you must put them in coordinate points or parenthesis (x,y).\[x~intercept = (2,0)\]\[y~intercept = (0,-\frac{ 2 }{ 3 })\]

OpenStudy (salwan):

thanks alot

OpenStudy (anonymous):

You could have just use an online graphing website @salwan

OpenStudy (anonymous):

No, the problem with online graphing websites is that they don't show work at all. Teachers love the process in between not the answer. After all it's putting down what you're supposed to know from class.

OpenStudy (anonymous):

@SCARLET_SPIDER